OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 3x + 5y = 3 x + 2y = 0

OpenStudy (anonymous):

First solve for x in the 2nd equation by subtracting 2y on both sides of the equation x + 2y = 0 x = -2y Now substitute what x equals in the 1st equation 3x + 5y = 3 3(-2y) + 5y = 3 -6y + 5y = 3 -y = 3 y = -3 Now that we know what y is, we can find out what x equals x + 2y = 0 x + 2(-3) = 0 x - 6 = 0 x = 6

OpenStudy (anonymous):

lol you got that straight off of yahoo xD